Operations Report – April 22nd

rew 211 Operations Report 22-APR-2019

SOL: 1

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 11:30

On (last night): 7:30 pm

Off (this morning): 7 am

Solar— SOC 88 % at 6:45 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~420 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 115 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 6 units

Toilet tank emptied: yes

Deimos rover used: no (rover is out of order according to Director)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 88.8h

Beginning charge: 100

Ending charge: 50

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: yes

Hours: 89.8h

Beginning charge: 100

Ending charge: 83

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: We cleaned all helmets and boiled earphones. A new plastic box was cleaned too and now contains all suit-related supplies.

Summary of Hab operations: Floor of the toilets and the whole lower deck were cleaned, and the mess that was lying on the desk was sorted.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Two operations were done: building stands with wood and the handsaw for the UWB anchors by Simon, and Carl assembled its probes with clamps.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Someone slightly twisted a bit his ankle while climbing down a ladder on the upper deck. Not really a problem.

Questions, concerns and requests to Mission Support:

* The hot water heater does not seem to work. Water does flow into the hot water pipes and taps, but keeps being cold.

Perhaps a problem with the propane supply ?

* The pump that flows from the loft tank to the taps seems to be leaking. The cardboard on the ceiling of the rightmost room is getting wet.

It looks that at least one previous crew was aware as we found a few sponges placed under the pump.

We fear that some long-term damage could happen to the cardboard and the wood that is immediately underneath it.

Additionally there is electric wiring up there, and water could enter in contact with it.

* I doubt that we will have enough water for the whole mission. We are trying to limit our consumption, but as a crew of eight even drinking water represents quite a cost.

Journalist Report – April 22nd

Today was our first simulation day. We woke up at 7:30 am and we had breakfast at 8:00. We ate cornflakes with dehydrated milk and orange juice. A part of crew started the day by gym while others did the housework. The first floor is now clean.

At 9 am, each of us worked on his experience (read the scientist’s report for more information).

At 10 am, Maxime and Eléonore tried to use solar telescope but were only able to use it in manual mode. After that, Eléonore and Simon prepared the lunch. We ate chiliconcarn with pasta. It was delicious. After meal, we had free time.

At 2pm, we continued our experiences. Each member find a place to work. At 4 pm, Carl, Eleonore, Simon and Chloé went to EVA to explore Mars. Chloé took soil samples for her own experiment.

At 5:15 pm, someone was on ladder when he twisted his ankle. But he seems better now.

At 6 pm, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. After the Cap Com, we will try to make sushi with basil from the greenhouse. At the end of the day, we will do a crew meeting to speak about the day.

Science Report – April 22nd

Crew 211 – UCL to Mars

Submitted by Science Officer Maxime Bernard

Commander Carl-Henrik Dahlqvist

Our commandant started to assemble the structure of his CubeSat that he printed prior to his arrival on Mars.

XO Officer

Simon started to build his wood stands on which he is going to fix his antennas.

HSO Officer Benjamin Flasse

This morning Benjamin gathered the data from the sleep monitoring he did on Eleonore.

Right after that, he measured everybody’s weight, tension and brain awakeness.

In addition to that, he took measures like fat density,bone quality and water repartition intra/extra cellular using Biody Xpert to monitor our body evolution along our stay here.

He spent the rest of the day making the inventory of what was available in the first aid kits. Most of the compresses were outdated and others were opened. The purchase of new medical supplies should be prioritized and more info will be available in one of the upcoming HSO reports.

Crew Astronomer Eleonore Lieffrig

Our astronomer was able to use the solar telescope for the first time today in manual mode.

Crew Engineer Julien Amalaberque

He is checking the parameters of an Intel RealSense D435i depth camera in order to evaluate which amount of stream data the positioning algorithm will be able to handle without being overwhelmed.

Crew Biologist Chloé Peduzzi

She spent the morning one the science dome preparing her culture medium in which she adds strains of spirulina.

She then disposed them in the greenhouse under two different light conditions before realizing it was way too hot in there. Therefore, she instead disposed them into the Science Dome. She encountered another problem we she realized she was sold the wrong transformer, 240V to 120V instead of 120V to 240V.

She went on EVA in the afternoon with three other crew members to gather soil samples.

Crew Journalist Nathan Pechon

Today, Nathan started to ask us questions about what we thought would be a good juridical baseline of a martian colony.

He then worked on the data he gathered during our talks to work on his first law article about what type of political regime should be adopted on Mars

Crew Scientist Maxime Bernard

Still waiting for the delivery of his muon detector that should have arrived on the 19th of April… The parcel seems to be somewhere between Salt Lake City and Hanksville.

Greenhab Report – April 22nd

Crew 211 GreenHab Report – 22-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 2

Shade cloth: 80%

Average temperatures:

Low: 13°C

High: 42°C

Hours of supplemental light: 0

Daily water usage for crops: 2 gallons + 2 gallons

Water in Blue Tank: 115 gallons

Time(s) of watering for crops: 9:00 + 18:00

Changes to crops: Still hot today. Some spinaches were stressed by the heat, but all the other plants still look healthy. 5 flowers are visible on the tomato plant and there are some flowers on Basil too. Romaine lettuce grow well.

Narrative: Still hot today despite the fan was turned up to 2. We have harvested 17,6g of Basil and 12g of Radish for the dinner.

Harvest: 17,6g Basil + 12g Radish

Support/supplies needed: None

Astronomy Report – April 22nd

Astronomy Report
Name: Lieffrig Eléonore

Crew: 211
Date: 22 April 2019

MDRS ROBOTIC OBSERVATORY

I’m planning to choose the quasars that I will observe and give the instructions to the telescope by tomorrow afternoon.

Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: None

Images submitted with this report: None

Problems Encountered: None

MUSK OBSERVATORY

Solar Features Observed: Edges of the sun, but no picture were taken because I had to go on EVA.

Problems Encountered: None

EVA Report #1 – April 22nd

EVA #1

Author: Eléonore Lieffrig

Purpose of EVA: Exploration and collection of soil samples

Start time: 4:30 pm

End time: 6:00 pm

Narrative:

We have had our first EVA this afternoon which lasted for two hours. We travelled to the hills on the left of the Cow Dung Road between the Robert’S Rock Garden and Zurbin’s head to take a few soil samples to be used by Chloé, the GreenHab officer, for one of her two experiments. The EVA went very well and the sunny weather allowed us to take some beautiful pictures of the Martian environment surrounding the MDRS.

Destination: Latitude: 4248000, longitude: 519000 (UTM NAD27 CONUS)

EVA Participants: Carl-Henrik Dahlqvist (commander of the EVA), Simon Collignon (medical officer), Chloé Peduzzi, Eléonore Lieffrig

Roads and routes per MDRS Map: Cow Dung Road 0110

Mode of travel: Driving and hiking

Sol Summary – April 22nd

Crew 211 Sol 1 Summary Report 22-APR-2019

Sol: 1

Summary Title: First day on Mars: first experiments and EVA

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: We start the day with a fitness session of 15 minutes which definitely helped some crew members to wake up. After the breakfast, we spent one hour and a half to clean the Hab. Most of the crew members have then started their respective experiments (see crew scientist report) while the crew astronomer started to work on the solar telescope. After the lunch, we all worked on our experiments until 4 pm when half of the crew went in EVA. As the UCL to Mars crew likes challenges, we will try tonight to prepare the first Martian Sushis and end the day with a meeting to discuss about today’s achievements and tomorrow’s objectives.

Look Ahead Plan: An EVA is planned in the morning to explore the Candor Chasma. Half of the crew will stay at the base to work on their experiments. Depending on the weather, the crew astronomer will try to get the first pictures of the sun and observed some fainter structures as the visual observations didn’t reveal any large perturbations in the solar atmosphere. The afternoon will again be devoted to the experiments.

Anomalies in work: Water tank pump leaking, water heater and heater in the Hab lab not working

Weather: Hot, sunny and fresh breeze.

Crew Physical Status: Good except for one of the crew member who twisted his ankle.

EVA: We have had our first EVA this afternoon which lasted for two hours. We traveled to the hills on the left of the Cow Dung Road between the Robert’S Rock Garden and Zurbin’s head to take a few soil samples to be used by Chloé, the GreenHab officer, for one of her two experiments. The EVA went very well and the sunny weather allowed us to take some beautiful pictures of the Martian environment surrounding the MDRS.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Operations Report,Astronomy Report, Scientist Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None